Which ordered pair is a solution of the system?
x+y =5
3x+4y =14
(1,4),(6,-1)or(3,2)

x+2y =7
3x-7y =8
(-1,4),(5,1)or(3,2)

x-5y =-2
2x-3y =10
(-2,0),(3,2)or(8,2)

Answer by elima(1433)

For each one is all you have to do is plug in the x and y given to you, and find which pair solves BOTH equations.
x+y =5
3x+4y =14
(1,4),(6,-1)or(3,2)
(1)+(4)=5
This solves this equation, lets check the other;
3(1)+4(4)=14
3+16=14
No, it does not solve this one, so the ordered pair does not work.
Let's try the second;
6+-1=5
Yes,
3(6)+4(-1)=14
18-4=14
Yes, so this ordered pair works; now try the last;
3+2=5
3(3)+4(2)=14
9+8=14
No, so the only one is;
(6,-1)
Now you try the others;
=)